OpenStudy (anonymous):

Let f(x) = 7x - 13. Find f-1(x). I'll medal

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (dangerousjesse):

Take the integral: integral \[\int\limits(7 x-13) dx \] Integrate the sum term by term and factor out constants: \[ = 7 \int\limits x d x-13 \int\limits 1 d x \] The integral of x is x^2/2: \[= (7 x^2)/2-13 \int\limits 1 dx \] The integral of 1 is x: Answer: \[ = (7 x^2)/2-13 x+constant \]

OpenStudy (dangerousjesse):

Ignore that first random "integral"

OpenStudy (anonymous):

Integration...why?

OpenStudy (anonymous):

He's looking for inverse, not area.

OpenStudy (dangerousjesse):

Because that was the most simple way of answering this question.. Holy crap. I completely misunderstood this question.

OpenStudy (anonymous):

Hahahaaaa

OpenStudy (dangerousjesse):

I wasted so much time on that -.-

OpenStudy (anonymous):

\[f(x) = 7x-13 ~~~ \implies y = 7x-13\] \[x = 7y-13 \] Solve for y now \[y = \frac{ x+13 }{ 7 } \implies f ^{-1}(x) = \frac{ x+13 }{ 7 }\]